OpenStudy (abbycross167):

How would I find the equation for the points (1,31) and (3,34)? can someone please help me?

TheSmartOne (thesmartone):

First find the slope :)

OpenStudy (alexandervonhumboldt2):

use the magic slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (alexandervonhumboldt2):

substitute your points and solve

OpenStudy (abbycross167):

m=34/3-31/1 would that work?

OpenStudy (alexandervonhumboldt2):

\[\frac{ 31-34 }{ 1-3 }=\frac{ -3 }{ -2 }=\frac{ 3 }{ 2 }\]

OpenStudy (alexandervonhumboldt2):

now we have y=mx+b where m is slope and b is y-intercept: m=3/2 subtitute one of the coordinates into the equation and solv for b: 31=3/2*1+b b=29.5 thus the equation is y=3/2x+29.5

OpenStudy (abbycross167):

oh thank you very much!!